Solve the system by substitution. x + y = 12
y = 2x

Answer by tutorcecilia(2144)

x + y = 12
y = 2x
.
x+(2x)=12 [Substitute (y=2x) for y]
3x=12
x=4
.
4+y=12 [solve for y]
y=12-4
y=8
.
checking:
x+y=12
(4)+(8)=12
12=12 [checks out]
and
y=2x
8=2(4)
8=8 [also checks out]